# Geometric Uncertainty for Detecting and Correcting Hallucinations in LLMs

### v1

@@ 5 Results
We evaluate our geometric framework for hallucination detection across multiple benchmarks and models. For global uncertainty, we test the Geometric Volume’s ability to classify response sets as reliable or hallucinated. For local uncertainty, we assess the Geometric Suspicion score in reducing hallucinations via a Best-of-N selection strategy. Each experiment is repeated three times to mitigate sampling stochasticity, and we report the mean and standard deviation of all metrics.
@@ 5.2 Reducing Hallucination Rates with Local Uncertainty
We show in Table 2 that using our local uncertainty metrics, we are able to reduce hallucination rates across all models and datasets investigated. Our method is particularly effective in ‘mid-hallucination’ subsets (described further in Section 6), which would be the primary intended use case for our hallucination reduction method, deployed in conjunction with a global uncertainty method for hallucination detection.
6 Discussion
Our local uncertainty framework has varying efficacy depending on the context in which it is applied. Let us consider three different situations of sampling-based uncertainty analysis, where for a given input we sample multiple times with a non-zero temperature.
Firstly, a ‘high-hallucination’ case, where the majority of sampled answers are in fact hallucinations. In this case, we would expect a high semantic dispersion, resulting in a high global uncertainty score. In such a case, particularly for safety-critical applications such as healthcare, we might decide that the model has insufficient capacity to answer the question, and escalate or defer the issue at hand. Secondly, a ‘low-hallucination’ case, where a minority of sampled answers are hallucinations. This should result in low semantic dispersion and low global uncertainty score, in which case we could propagate the original, zero-temperature response. Finally, the ‘mid-hallucination’ case, where there is no clear majority of hallucinations or non-hallucinations, and global uncertainty analysis yields a middling score. In these cases the model may have sufficient capacity to answer the question well, but we might not be confident in the original, zero-temperature response. These are the crucial cases where we can leverage local uncertainty scores to reduce hallucination rates.
(a) A low rank for each component
@@ 6 Discussion
Figure 3: t-SNE plots for cases where we were able to prevent hallucinations with use of our local uncertainty metrics, using the TriviaQA dataset and GPT-4o-mini model. In each plot, the red X shows the original model answer, sampled with a temperature of zero, which was determined to be a hallucination by a judge LLM. The other points show the batch of $n=20$ answers sampled from the same model with a temperature of one. The blue cross shows the answer selected by our framework (i.e. with lowest suspicion), which was determined by the same judge LLM to be a non-hallucination.
In Figure 3, we present t-SNE plots of cases where our local uncertainty framework was able to ‘flip’ a model response from a hallucination to a non-hallucination. The different examples show how the component terms of Geometric Suspicion contribute to selecting a desirable sample from the batch.
@@
Figure 3(a) shows a straightforward example where all three of the component metrics rank highly. The correct answer is in a dense local cluster, near a global consensus point, and can be reconstructed using commonly used archetypes. Figure 3(b) shows a slightly harder example, where there are multiple clusters of semantically similar answers (all are countries or some slight variation thereof). Here the desired answer is not within the densest local cluster or the closest to a global consensus, but has lower suspicion due to its non-reliance on rarely used archetypes. The example in Figure 3(c) can be used to understand the necessity of both local density and distance from consensus metrics. Here we have a few dense clusters, resulting in high local density for the majority of the batch. The global semantic consensus point however is closer to the ‘Indonesia’ cluster than it is the ‘Not specified’ or ‘Unknown entity’ clusters. This is because the hallucinated answers of ‘Brazil’ or ‘Pakistan’ are countries and semantically closer to the correct answer, moving the global consensus further in this direction. Finally, Figure 3(d) shows our framework can be useful in cases of diverse answer batches without dense local clusters. In this example approximately half of the answers are semantically unique, but the use of commonly used archetypes in the ‘W.H. Auden’ cluster allows us to find a low suspicion example which proves to be correct.
@@ Appendix A Theoretical Analysis
A.1 Simplicial Volume as a Proxy for Uncertainty
We now formalize this relationship by proving that the volume of the archetypal simplex provides an upper bound on the differential entropy of any distribution supported within it, thereby linking geometric dispersion to semantic uncertainty.
Theorem 1.
Proof.
See Appendix  B. ∎
@@ Appendix B Proof of Theorem 1
Appendix C Experimental Setup
C.1 Benchmarks and Baselines
@@
To evaluate the detection of external hallucinations arising from ambiguous user prompts, we use the CLAMBER benchmark (Zhang et al., 2024). We use 3,202 queries from the dataset, where the model’s uncertainty should reflect the ambiguity of the input rather than a lack of internal knowledge.
@@
We also evaluate on TriviaQA, a short-form question answering benchmark (Joshi et al., 2017). From the TriviaQA dataset, we construct a balanced test set of 1,000 samples, consisting of 500 hallucinated and 500 non-hallucinated examples.
@@
We also used ScienceQA as a benchmark (Lu et al., 2022), which is a multimodal multiple-choice dataset designed to assess scientific reasoning across natural, social, and language sciences. For our evaluation, we filtered the test set to include only text-only questions and created a balanced benchmark of 200 hallucinating and 200 non-hallucinating examples for each LLM.
@@
To assess performance in a critical, high-stakes domain, we use a medical question-answering benchmark subset (500 examples) from MedQA and MedMcQA (Zhang et al., 2023). This dataset tests the model’s ability to provide factually correct answers to medical questions, where hallucinations carry significant risk. MedicalQA is a long-form question answering benchmark, unlike TriviaQA, and provides deeper insight into hallucination in real-world applications.
@@
Finally, to evaluate uncertainty in real-world medical scenarios, we leverage 201 real patient questions from the K-QA dataset (Manes et al., 2024), a long-form generation benchmark with expert-written answers reviewed by licensed physicians. These detailed responses serve as high-quality references for assessing the accuracy and reliability of model outputs.
C.2 Dataset Curation
@@
For the CLAMBER benchmark, we evaluate each of the 3,202 queries, where each $q_{i}$ is annotated with a binary label $y_{i}\in{0,1}$ indicating whether it is ambiguous. We generate $n=20$ perturbations $\{q_{i}^{(j)}\}_{j=1}^{n}$ for each query by prompting each LLM following previous work (Li et al., 2025), and obtain embeddings $\{x_{i}^{(j)}\}_{j=1}^{n}$ via the Alibaba-NLP/gte-Qwen2-1.5B-instruct model.
@@
To create hallucination detection datasets from traditional QA benchmarks, we convert question–answer pairs $(q_{i},a_{i}^{\text{ref}})$} into labeled examples. For each question $q_{i}$, we generate a response $r_{i}=\text{LLM}(q_{i})$ using LLM. We then assign a hallucination label $y_{i}\in\{0,1\}$ by comparing $r_{i}$ to the reference answer $a_{i}^{\text{ref}}$. In high-stakes domains such as medical QA, this comparison is performed by an LLM-as-a-judge (Gu et al., 2024) (GPT-4o) using expert-annotated references. For example, given a real-world clinical question and its verified answer, we label $r_{i}$ as hallucinated if the LLM judge determines it deviates from the reference. For datasets like TriviaQA, where answers are typically short (1–3 words), we instead compute ROUGE-L F1 and label $r_{i}$ as hallucinated if $\text{ROUGE}(r_{i},a_{i}^{\text{ref}})<0.3$, following previous work (Li et al., 2025). When initially answering questions, we set the generation temperature to 0. Later, for sampling perturbed responses, we use a temperature of 1.0 to induce diversity. Finally, for each $q_{i}$, we sample $n=20$ perturbed responses $\{r_{i}^{(j)}\}_{j=1}^{n}$ and embed them via $\{\mathbf{x_{i}^{(j)}}=E(r_{i}^{(j)})\}_{j=1}^{n}$, using the same embedding pipeline as in the external uncertainty setting.
C.3 Implementation Details
@@
We set the PCA dimension to 15 and number of archetypes $K=16$ for each benchmark. The Archetypal Analysis optimization is run for 2000 steps on each batch of sampled responses. For a fair comparison, both our method and the Semantic Volume baseline require an uncertainty threshold $\tau$ to make a binary prediction (hallucination vs. not) (Li et al., 2025). Following prior work, we tune this threshold on a held-out validation split (10% of the data), selecting the $\tau$ that maximizes the F1-score. Local uncertainty experiments are performed with $k=5$ nearest neighbours for the local density metric.
@@ Appendix D Selection of Geometric Suspicion Terms
For each model $M$ and dataset $D$ investigated, we have a set of $N_{D,M}$ questions. For each question, we have a ‘default’ answer, sampled with a temperature $T=0$, and a corresponding hallucination label as determined by a judge LLM (as described in Appendix Section C). For each question, we have in addition $n=20$ answers sampled with temperature $T=1$, for which we generate hallucination labels as for the default answer. For each batch of answers, we then have the objects derived from our archetypal analysis: the answer embeddings $\mathbf{X}$, the convex decomposition coefficients $\mathbf{A}$, and the archetype embeddings $\mathbf{Z}$.
@@
To test these, we compiled several terms of interest and compared them across our datasets. To measure the density of points within the embedding space, we used Local Density with $k=5$ (as described in Section 3.3) and Voronoi cell volume. The Voronoi cell of an embedding $\mathbf{x}_{i}$ is the region of space closer to $\mathbf{x}_{i}$ than to any other point in the batch. A larger cell volume implies the response lies in a sparser region. As direct computation of these volumes in high-dimensional space is intractable, we approximate them via the dual Delaunay triangulation. First, for each batch of embeddings $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$, we reduce their dimensionality to a low-dimensional space (typically 3D) using Principal Component Analysis (PCA). We then compute the Delaunay triangulation of these reduced points. The volume of a response’s Voronoi cell is approximated by summing the volumes of all Delaunay simplices that have the response’s embedding as a vertex. A larger estimated volume signifies that a response lies in a sparser region of the local embedding space and therefore might be considered more suspicious.
@@
To assess the archetypal composition, we used Usage Rarity, as described in Section 3.3 and Geometric Entropy. The latter is the Shannon entropy (Shannon, 1948) of the archetypal reconstruction coefficients. The $i$-th row of $\mathbf{A}$, denoted $\bm{\alpha}_{i}=[A_{i1},\dots,A_{iK}]$, is a probability distribution describing how the pre-processed response embedding $\mathbf{x}_{i}$ is reconstructed from the set of learned archetypes $\mathbf{Z}$. A response that aligns clearly with a single archetype will have a sparse, low-entropy $\bm{\alpha}_{i}$, while a response that lies between multiple archetypes will have a more uniform, high-entropy $\bm{\alpha}_{i}$.

### v2

@@ 5 Results
We evaluate our geometric framework for hallucination detection across multiple benchmarks and models. For global uncertainty, we test Geometric Volume’s ability to classify $r_{\text{default}}$ as reliable or hallucinated given a response set. For local uncertainty, we assess the Geometric Suspicion score in reducing hallucinations via a Best-of-N selection strategy. Each experiment is repeated three times to mitigate sampling stochasticity, and we report the mean and standard deviation of all metrics.
@@ 5.2 Reducing Hallucination Rates with Local Uncertainty
We show in Table 2 that using our local uncertainty metric, we are able to reduce hallucination rates across all models and datasets investigated. In general it surpasses baselines across models and datasets, with particularly dominant performance in the MedicalQA task.
We attribute this performance primarily due to the combination of terms employed, compared to baselines which use only one term. In Figure 3 we show how reliance on one notion of ‘confidence’, such as local density or distance from consensus, would lead to Best-of-N selection of a hallucinated response, whereas our composite metric selects a correct response.
(a) A high p-score for each component
@@
Figure 3: UMAP plots for cases where we were able to correct hallucinations with our local uncertainty metric, using the TriviaQA dataset and GPT-4o-mini model. In each plot, the red X shows $r_{\text{default}}$, which was determined to be a hallucination by a judge LLM. The other points show the batch of $n=20$ answers sampled from the same model with a temperature of one. The blue cross shows the answer selected by our framework (i.e. with lowest suspicion), which was determined by the same judge LLM to be a non-hallucination.
@@
Figure 3(a) shows a straightforward example where all three of the component terms have high p-values. The correct answer is in a dense local cluster, near a global consensus point, and can be reconstructed using commonly used archetypes. Figure 3(b) shows a slightly harder example, where there are multiple clusters of semantically similar answers. Here the desired answer is not within the densest local cluster or the closest to a global consensus, but has lower suspicion due to its non-reliance on rarely used archetypes. The example in Figure 3(c) can be used to understand the necessity of both local density and distance from consensus metrics. Here we have a few dense clusters, resulting in high local density for the majority of the batch. The global semantic consensus point however is closer to the ‘Indonesia’ cluster than it is the ‘Not specified’ or ‘Unknown entity’ clusters. This is because the hallucinated answers of ‘Brazil’ or ‘Pakistan’ are countries and semantically closer to the correct answer, moving the global consensus further in this direction. Finally, Figure 3(d) shows our framework can be useful in cases of diverse answer batches without dense local clusters. In this example approximately half of the answers are semantically unique, but the use of commonly used archetypes in the ‘Michael Nesmith’ cluster allows us to find a low suspicion example which proves to be correct.
An interesting case is Qwen3-8B for the TriviaQA dataset, where all methods tested fail to reduce the hallucination rate. We find that compared to other model/dataset combinations, in this case the model is more often confidently wrong when it hallucinates: that is, very few of the sampled answers are non-hallucinations, so choosing an optimal answer becomes a difficult task. We comment on this method limitation in Section 6 and provide the evidence for this conclusion in Appendix Section E.
6 Discussion
Our geometric framework views language model UQ as modelling a probability distribution over answers in semantic space. The entropy of this distribution, bounded by geometric volume, yields useful global uncertainty signals. Like previous methods such as Semantic Density, we then frame local UQ as finding optimal points within the distribution. Our convex hull approach is a novel method for defining the semantic distribution boundaries, whilst our local metric approximates a ‘confidence’ distribution within the boundaries without access to model likelihoods.
Our framework has several limitations, which we aim to address in future work. Firstly, it relies on the capability of a single embedding vector to sufficiently capture the semantic content of an answer. As answers get longer and semantically more complex, this assumption may break down. Recent work has investigated this and suggests that the principle of semantic distribution analysis with a single embedding vector holds promise for answers of up to 1000 words (bhardwaj2025embeddingtrustsemanticisotropy), but it remains to be proven for complex datasets beyond QA tasks.
Our method relies on the principle of semantic dispersion in sampled answers predicting hallucination. It is therefore most effective in settings where the model is uncertain and wrong, i.e. there is sufficient diversity in the sampled answers. There exist cases where the model is confident and wrong (no diversity in sampled answers). White-box methods such as factuality probes (han-etal-2025-simple) may be able to address these cases, but to our knowledge no black-box methods have successfully solved this problem.
@@ A.1 Proof of Theorem 1
Appendix B Experimental Setup
B.1 Benchmarks and Baselines
@@
To evaluate the detection of external hallucinations arising from ambiguous user prompts, we use the CLAMBER benchmark (zhang2024clamber). We use 3,202 queries from the dataset, where the model’s uncertainty should reflect the ambiguity of the input rather than a lack of internal knowledge.
@@
We also evaluate on TriviaQA, a short-form question answering benchmark (joshi2017triviaqa). From the TriviaQA dataset, we construct a balanced test set of 1,000 samples, consisting of 500 hallucinated and 500 non-hallucinated examples.
@@
We also used ScienceQA as a benchmark (lu2022learn), which is a multimodal multiple-choice dataset designed to assess scientific reasoning across natural, social, and language sciences. For our evaluation, we filtered the test set to include only text-only questions and created a balanced benchmark of 200 hallucinating and 200 non-hallucinating examples for each LLM.
@@
To assess performance in a critical, high-stakes domain, we use a medical question-answering benchmark subset (500 examples) from MedQA and MedMcQA (zhang2023huatuogpt). This dataset tests the model’s ability to provide factually correct answers to medical questions, where hallucinations carry significant risk. MedicalQA is a long-form question answering benchmark, unlike TriviaQA, and provides deeper insight into hallucination in real-world applications.
@@
Finally, to evaluate uncertainty in real-world medical scenarios, we leverage 201 real patient questions from the K-QA dataset (manes2024k), a long-form generation benchmark with expert-written answers reviewed by licensed physicians. These detailed responses serve as high-quality references for assessing the accuracy and reliability of model outputs.
B.2 Dataset Curation
@@
For the CLAMBER benchmark, we evaluate each of the 3,202 queries, where each $q_{i}$ is annotated with a binary label $y_{i}\in{0,1}$ indicating whether it is ambiguous. We generate $n=20$ perturbations $\{q_{i}^{(j)}\}_{j=1}^{n}$ for each query by prompting each LLM following previous work (li2025semantic), and obtain embeddings $\{x_{i}^{(j)}\}_{j=1}^{n}$ via the Alibaba-NLP/gte-Qwen2-1.5B-instruct model.
@@
To create hallucination detection datasets from traditional QA benchmarks, we convert question–answer pairs $(q_{i},a_{i}^{\text{ref}})$} into labeled examples. For each question $q_{i}$, we generate a response $r_{i}=\text{LLM}(q_{i})$ using LLM. We then assign a hallucination label $y_{i}\in\{0,1\}$ by comparing $r_{i}$ to the reference answer $a_{i}^{\text{ref}}$. In high-stakes domains such as medical QA, this comparison is performed by an LLM-as-a-judge (gu2024survey) (GPT-4o) using expert-annotated references. For example, given a real-world clinical question and its verified answer, we label $r_{i}$ as hallucinated if the LLM judge determines it deviates from the reference. For datasets like TriviaQA, where answers are typically short (1–3 words), we instead compute ROUGE-L F1 and label $r_{i}$ as hallucinated if $\text{ROUGE}(r_{i},a_{i}^{\text{ref}})<0.3$, following previous work (li2025semantic). When initially answering questions, we set the generation temperature to 0. Later, for sampling perturbed responses, we use a temperature of 1.0 to induce diversity. Finally, for each $q_{i}$, we sample $n=20$ perturbed responses $\{r_{i}^{(j)}\}_{j=1}^{n}$ and embed them via $\{\mathbf{x_{i}^{(j)}}=E(r_{i}^{(j)})\}_{j=1}^{n}$, using the same embedding pipeline as in the external uncertainty setting.
B.3 Implementation Details
@@
We set the PCA dimension to 15 and number of archetypes $K=16$ for each benchmark. The Archetypal Analysis optimization is run for 2000 steps on each batch of sampled responses. For a fair comparison, both our method and the Semantic Volume baseline require an uncertainty threshold $\tau$ to make a binary prediction (hallucination vs. not) (li2025semantic). Following prior work, we tune this threshold on a held-out validation split (10% of the data), selecting the $\tau$ that maximizes the F1-score. Local uncertainty experiments are performed with $k=5$ nearest neighbours for the local density metric.
Baselines
For Eccentricity and Degree (lin2024generatingconfidenceuncertaintyquantification), we adapt the code from previous work (https://github.com/zlin7/UQ-NLG) and use the entailment variant with the Deberta-Large model (https://huggingface.co/microsoft/deberta-large and an eigenvalue threshold of 0.7.
Appendix C Ablations
C.1 Local Uncertainty
We probe robustness of our local uncertainty estimator along three axes: number of sampled answers, PCA dimension, and number of archetypes.
First, Table 3 varies the number of sampled answers $n$. We maintain our fixed answer pool of 20 answers from main experiments: subsets of size $n$ are sampled uniformly without replacement from each question’s fixed answer pool. When $n<n_{\text{all}}$ we evaluate $R=30$ repeated subsets per question. We fit PCA on each subset, so the effective dimensionality obeys $d^{\prime}\!\leq\!n-1$. We scale the archetype count as $K=\mathrm{round}(\rho n)$ with $\rho=K_{\mathrm{base}}/n_{\mathrm{base}}$ (16 / 20 for our chosen settings), then clip K to satisfy the assumptions of our theorem: $K\leq n$ and $K\leq d^{\prime}+1$, ensuring a non-degenerate, identifiable simplex in $d^{\prime}$ dimensions. As expected, the hallucination correction rate increases with higher $n$, as the sampled set becomes more likely to contain one or more correct answers.
Second, Tables 4 and 5 vary the PCA dimension $d^{\prime}$ at fixed $n=20$: we show a fixed-$K$ view ($K=3$) and a capacity-matched view ($K=\min(d^{\prime}+1,K_{\mathrm{base}},n)$) that respects the simplex capacity of a $d^{\prime}$-dimensional space. In both cases, increasing $d^{\prime}$ up to our chosen setting improves the hallucination detection performance. We suggest this is due to low PCA dimensions being unable to capture sufficient geometric details of the semantic answer space.
Finally, Table 6 sweeps the number of archetypes $K$ at fixed $d^{\prime}$ (default $d^{\prime}=15$) while enforcing the same feasibility limits. Interestingly our local metric is able to perform well even with very low numbers of archetypes; we suggest this is due to two of the three component terms using the raw answer embeddings, such that altering the ‘usage rarity’ term does not drastically affect results.
Statistics.
[…]
$n$ AUROC AUARC $\Delta H$
5 $0.540\,\pm\,0.019$ $576\,\pm\,0.017 0.0$120.012
10 0.550± $50\,\pm\,0.012$ $593\,\pm\,0.015 0.0$310.031
15 $0.547\,\pm\,0.008$ $587\,\pm\,0.019 0.0$390.039
20 $0.556\,\pm\,0.007$ $585\,\pm\,0.019 0.0$680.068
Table 4: Ablation over PCA dimension $d^{\prime}$ at fixed $n=20$. We fit a single PCA basis on all 20 answers and compare by truncation. We fix $K=3$ so results remain feasible down to $d^{\prime}=2$.
$d^{\prime}$ AUROC AUARC $\Delta H$
2 $0.518\,\pm\,0.005$ $564\,\pm\,0.017 0.0$21-0.021
3 $0.521\,\pm\,0.010$ $567\,\pm\,0.015 0.0$22-0.022
5 $0.535\,\pm\,0.003$ $575\,\pm\,0.017 0.0$440.044
10 $0.551\,\pm\,0.008$ $581\,\pm\,0.019 0.0$320.032
15 $0.554\,\pm\,0.003$ $585\,\pm\,0.018 0.0$620.062
Table 5: Ablation over PCA dimension $d^{\prime}$ at fixed $n=20$ with capacity-matched archetypes: $K=\min(d^{\prime}+1,K_{\mathrm{base}},n)$. This reflects the geometric limit that at most $d^{\prime}+1$ archetypes can be supported in a $d^{\prime}$-dimensional space.
$d^{\prime}$ AUROC AUARC $\Delta H$
2 $0.519\,\pm\,0.003$ $565\,\pm\,0.019 0.0$18-0.018
3 $0.520\,\pm\,0.007$ $568\,\pm\,0.016 0.0$160.016
5 $0.534\,\pm\,0.003$ $573\,\pm\,0.020 0.0$250.025
10 $0.552\,\pm\,0.005$ $583\,\pm\,0.018 0.0$520.052
15 $0.559\,\pm\,0.004$ $586\,\pm\,0.021 0.0$680.068
Table 6: Ablation over the number of archetypes $K$ at fixed $n=20$ and fixed $d^{\prime}$ (default $d^{\prime}=15$). We enforce feasibility $K\leq n$ and $K\leq d^{\prime}+1$.
$K$ AUROC AUARC $\Delta H$
[…]
8 $0.556\,\pm\,0.003$ $585\,\pm\,0.019 0.0$630.063
12 $0.556\,\pm\,0.003$ $586\,\pm\,0.020 0.0$540.054
16 $0.556\,\pm\,0.003$ $585\,\pm\,0.020 0.0$640.064
C.2 Global Uncertainty
We report similar ablations for global uncertainty in Table 7, showing similarly that high $n$ and $d^{\prime}$ are desirable, with performance fairly constant as number of archetypes are varied. We choose $K=16$ in our main results for improved performance in our local metrics.
Table 7: Sensitivity Analysis on TriviaQA. We report Mean $\pm$ Std over 3 runs. (a) As expected, increasing the sample size leads to improved performance. (b) The method requires $d^{\prime}\geq 5$ to capture semantic structure; the poor performance at $d^{\prime}=2$ validates the necessity of high-dimensional volume over simple 2D area. (c) Geometric Volume is robust to the number of archetypes $K$, remaining stable even with a minimal simplex ($K=4$).
(a) Sample Size ($n$)
$n$ AUROC F1
5 0.705 $\pm$ 0.010 0.652 $\pm$ 0.017
10 0.731 $\pm$ 0.004 0.674 $\pm$ 0.012
15 0.744 $\pm$ 0.003 0.682 $\pm$ 0.014
20 0.752 $\pm$ 0.002 0.701 $\pm$ 0.004
(b) PCA Dim ($d^{\prime}$)
$d^{\prime}$ AUROC F1
2 0.692 $\pm$ 0.014 0.666 $\pm$ 0.002
5 0.746 $\pm$ 0.005 0.701 $\pm$ 0.007
10 0.752 $\pm$ 0.003 0.705 $\pm$ 0.004
15 0.753 $\pm$ 0.002 0.705 $\pm$ 0.002
(c) Archetypes ($K$)
$K$ AUROC F1
4 0.753 $\pm$ 0.001 0.706 $\pm$ 0.001
8 0.753 $\pm$ 0.002 0.704 $\pm$ 0.004
12 0.753 $\pm$ 0.002 0.707 $\pm$ 0.002
16 0.753 $\pm$ 0.002 0.705 $\pm$ 0.004
@@ Appendix D Selection of Geometric Suspicion Terms
For each model $M$ and dataset $D$ investigated, we have a set of $N_{D,M}$ questions. For each question, we have a ‘default’ answer, sampled with a temperature $T=0$, and a corresponding hallucination label as determined by a judge LLM (as described in Appendix Section B). For each question, we have in addition $n=20$ answers sampled with temperature $T=1$, for which we generate hallucination labels as for the default answer. For each batch of answers, we then have the objects derived from our archetypal analysis: the answer embeddings $\mathbf{X}$, the convex decomposition coefficients $\mathbf{A}$, and the archetype embeddings $\mathbf{Z}$.
@@
To test this, we compiled several terms of interest and compared them across our datasets. To measure the density of points within the embedding space, we used Local Density with $k=5$ (as described in Section 3.4) and Voronoi cell volume. The Voronoi cell of an embedding $\mathbf{x}_{i}$ is the region of space closer to $\mathbf{x}_{i}$ than to any other point in the batch. A larger cell volume implies the response lies in a sparser region. As direct computation of these volumes in high-dimensional space is intractable, we approximate them via the dual Delaunay triangulation. First, for each batch of embeddings $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$, we reduce their dimensionality to a low-dimensional space (typically 3D) using Principal Component Analysis (PCA). We then compute the Delaunay triangulation of these reduced points. The volume of a response’s Voronoi cell is approximated by summing the volumes of all Delaunay simplices that have the response’s embedding as a vertex. A larger estimated volume signifies that a response lies in a sparser region of the local embedding space and therefore might be considered more suspicious.
@@
To assess the archetypal composition, we used Usage Rarity, as described in Section 3.4 and Geometric Entropy. The latter is the Shannon entropy (shannon1948mathematical) of the archetypal reconstruction coefficients. The $i$-th row of $\mathbf{A}$, denoted $\bm{\alpha}_{i}=[A_{i1},\dots,A_{iK}]$, is a probability distribution describing how the pre-processed response embedding $\mathbf{x}_{i}$ is reconstructed from the set of learned archetypes $\mathbf{Z}$. A response that aligns clearly with a single archetype will have a sparse, low-entropy $\bm{\alpha}_{i}$, while a response that lies between multiple archetypes will have a more uniform, high-entropy $\bm{\alpha}_{i}$.
@@
Appendix E Local Uncertainty Efficacy
In Section 6, we note that the efficacy of local uncertainty measures inherently bounded by the quality of the sampled response set. To quantify this, we analyzed the relationship between the reduction in hallucination rate ($\Delta H$) and the proportion of correct answers available in the sampled batch for questions where the default response was a hallucination.
Figure 5 illustrates this relationship across all model and dataset combinations. The x-axis represents the Potential to Correct: defined as the median proportion of non-hallucinated answers found in the sampled batch given that the default greedy decode was incorrect. The y-axis shows the actual performance gain ($\Delta H$).
We observe a strong positive correlation. Notably, the case of Qwen3-8B on TriviaQA (the lowest performance in $\Delta H$ terms) demonstrates the ‘confidently wrong’ failure mode: while the model’s default answer is incorrect, the sampled batch also contains very few correct alternatives. In such regimes, the semantic space is collapsed around the hallucination, and no re-ranking method can be effective. Conversely, as the correctness proportion of the sampled batch improves (moving right), Geometric Suspicion successfully identifies and retrieves the correct answers.
Figure 5: Geometric Suspicion Utility vs. Corrective Potential. The reduction in hallucination rate ($\Delta H$) is plotted against the median proportion of correct answers found in the sampled batch for instances where the default answer was incorrect. The dashed curve represents a second-order polynomial regression fit highlighting the general scaling trend, while the shaded region indicates the 95% confidence interval. Performance naturally drops in ‘confidently wrong’ regimes where the sampled batch lacks correct alternatives, but improves as the corrective potential increases.
Appendix F Computational and Memory Complexity
The dominant cost of our framework arises from the archetypal analysis step, which alternates between coefficient and archetype updates. Each update involves matrix multiplications over $n$ sampled responses of reduced dimension $d^{\prime}$, resulting in a total time complexity of $\mathcal{O}(T\,nd^{\prime}K)$, where $T$ is the number of alternating optimization steps and $K$ the number of archetypes. In our setting, $d^{\prime}<n$ due to dimensionality reduction, and both $K$ and $T$ are small constants, making the method effectively linear in the number of responses and embedding dimensions.
The memory footprint is dominated by the response embeddings $\mathbf{X}\in\mathbb{R}^{n\times d^{\prime}}$ and coefficient matrices $\mathbf{A}\in\mathbb{R}^{K\times n}$ and $\mathbf{B}\in\mathbb{R}^{n\times K}$, yielding an overall space complexity of $\mathcal{O}(nd^{\prime}+nK+Kd^{\prime})$. Since $K\leq\min(n,d^{\prime}+1)$ in all experiments, the approach remains both computationally and memory efficient for large-scale response batches.